Tohban Report 2015-10-21
From RHESSI Wiki
|Start Date:||14 Oct 2015|
|End Date:||21 Sep 2015|
|List all reports|
Some activity, including a couple of M flares. Activity has been quieter the last couple of days:
How many GOES flares occurred?
Flares above B, C, M, X class were 16 51 4 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 7 32 3 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 5 16 2 0
There were RHESSI flares/GOES flares 142 / 71 over the time range 13-Oct-15 20-Oct-15
Decimation was active/vigorous for the full time period. Both front and rear detectors are off during SC night, except for 4 minutes on either side. The SSR is at 29%, with 6 contacts left for the day, and has been dropping to about 13% after each day's set of passes.
The cold plates, 1 and 2 were at 135.1 and 133.6 K, on 2015-10-19.
VC1 Gaps GAP START TIME GAP END TIME GAP (SEC)
2015-10-18T17:55:00.000 -- 2015-10-18T18:10:00.000 900.00000
Detectors 2 is unsegmented. Detector 4 is not recording events. Detector 9 is off. No threshold or high-voltage changes were made. Live time for each ok detector has increased throughout the week, including D8f, which has increased from ~0.45 live time to ~0.6. The reset rate on D8r is very high, but has dropped over the past week from 2.5e5 to 1.5e5.
On 14-Oct-2015, the attenuator thresholds were increased as discussed in that day's ops meeting. Background dead time was added to the original, 2008, settings to come up with the new thresholds. The most significant change was the thick IN (Max l1), which increased from 7% to ~13% dead time.
Both attenuators were out at the time of this commanding, 2015-287-22:42:52 UTC.
All of the threshold values, in dead time, can be found below. Please note, the l2 thresholds are included, but are not used in our current configuration. We only use the 0, 1, and 3 states.
Min l0 ; 0% Max l0 ; 11.0% (increase from 8.2%) Min l1 ; 7.45% Max l1 ; 12.9% (increase from 7.0%) Min l2 ; 6.6% (increase from 0.8%) Max l2 ; 20.8% (increase from 14.9%) Min l3 ; 7.84% Max l3 ; 100%
Attenuator logic seems to be working as expected for the thin attenuator. There were no thick attenuator motions this week.
For reference, the 2008 attenuator threshold values were:
Max l0 5.1% Min l1 1.5% Max l1 7.06% Min L3 2.35%
|HLAT Decimation||Rear decimation weight 6, no front decimation|
|Night time data (fronts)||+/- 4 minutes|
|Night time data (rears)||+/- 4 minutes|
|Require extra passes?||No|
|Requirement for moving pointer?||No|
|Attenuator operation||See above|
|Thin Att. in||11%|
|Thin Att. out||7.45%|
|Thick Att. in||12.9%|
|Thick Att. out||7.84%|
|Detector problems?||See notes above.|